# All Optical Echo State Network Reservoir Computing

### v2

@@ IV Results and Discussion
Figure 3: Comparison of the performance of the Sine-Square Classification test. Parameters: $N=10$, $gL=12$, $\mathrm{Pump/Input}=1500$
Good performance in machine learning hinges on the ability to approximate nonlinear input–output relationships and capture complex temporal dynamics. According to the Stone–Weierstrass theorem [31], the polynomial algebra property is central to achieving universality. Therefore, we selected three distinct tasks to evaluate these critical capabilities. They are the sine-square classification test, prediction of the NARMA10 system, and the Mackey-Glass time series prediction. Each task evaluates mainly the reservoir computer nonlinearity, polynomial algebra capabilities, and dynamic memory retention, respectively. Additionally, we address the nonlinear properties of the SBS interaction and, thus, the optimization of the activation function. To achieve good results in the tasks mentioned above, we require strong enough nonlinearity, while maintaining the convergence conditions of the ESN. We vary the value of the $gL$ product as well as the pump splitting ratio with these constraints in mind to optimize the performance of the ESN in the selected benchmark tests.
@@
For each test, we analyze the effect of the nonlinear strength of the SBS interaction, given by the $gL$ product value of the medium and the pump splitting ratio for the self-pumped nonlinear activation element, which refers to the ratio of the pump power to the input power. Here, $g$ represents the Brillouin gain and $L$ represents the interaction length. These values appear in Equation (7). The $gL$ product decides the nonlinear shape and the parking point of the activation function, as shown in Figure 1. The pump splitting ratio independently changes the parking point of the activation function, which allows us to optimize it for greater performance. It is important to note however that increasing the splitting ratio increases the power budget of the system. The parking point of the tunable activation function moves between -5 dBm to -10 dBm for the range of gL values that we have used for testing in this work, as shown in Figure 1. This means that, around the parking points, the pump power will be 100300100-300 mW for a splitting ratio of 1000. Since we have allowed for a maximum output $x_{k}$ value of $31$ dBm, the maximum pump power required in this scenario is $3.2$ W, which is quite a feasible power budget for the nonlinear activation elements. Better performance can generally be achieved for larger splitting ratios, which means that the best performance is restricted by the energy availability.
@@
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.
@@
In order to simulate the proposed architecture for the matrix-vector multiplication, the data at the input and output is always represented in dBm units, which allows access to negative elements as discussed in the previous section. However, the input dBm data is always converted to W units before the multiplication step. This must be done in order to accurately simulate the behavior of optical components such as VOAs and Brillouin amplifiers. The output is converted back to dBm for the detection phase, ensuring that the data is always represented in dBm units.
@@ IV.1 Sine Square Classification
This test is a time series classification task, commonly used in estimating the nonlinearity of computing reservoirs [32]. The input given is a series of randomly arranged sine and square waves, each 8 steps long. The task given to the ESN is to classify the current wave as sine or square. This test helps us understand the nonlinearity of the ESN. Distinguishing between smoothly varying sine waves and abruptly changing square waves requires mapping the data into a higher-dimensional hyperspace (i.e., nonlinear transformation), making classification easier. The results are shown in Figure 3.
On average, we see similar performance in terms of the NMSE with the software ESN using a sigmoid activation performing marginally better. We emphasize that the comparison between the software ESN and the all-optical ESN is not intended to show that one approach is superior to the other. Rather, the comparison serves as a validation check for the numerical simulations of our all-optical ESN. We anticipate that the performance of the software ESN and the all-optical ESN will be quite similar, as both implementations rely on the same underlying model, differing only slightly in their nonlinear activation functions. The reported values for this instance given in the figures are 0.14 for the software ESN and 0.12 for the optical ESN. These are median values from multiple computations.
@@
Figure 5: Comparison of the performance of the NARMA10 test. Parameters: $N=10$, $gL=12$, $\mathrm{Pump/Input}=1500$
The values reported by Slinkov et. al [22] allow the possibility of $gL$ values from 373-7 with fiber lengths of 363-6m at reasonable threshold powers of $1.26$ W which are very practical lengths of fiber. Experimental demonstrations have shown much higher levels of nonlinearity in the right materials [33, 28]. This shows promise for the translation of this architecture into an on-chip implementation, which has the benefit of improved scalability while maintaining the other advantages of this fiber-based approach.
@@ IV.2 NARMA10
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.
@@
Next, the nonlinear autoregressive moving average (NARMA) [34] task is used to test the ESN. This task mainly focuses on the polynomial algebra property of the all-optical ESN reservoir. The model for NARMA10, a version of NARMA popularly used for testing reservoir computers, is given by
@@
For this test, we used the following parameters of the all-optical ESN: $N=10,gL=12,$ and the pump/input ratio of 1500. On average we see marginally better performance from the software sigmoid ESN, but the NMSE values are comparable. We note that the performance difference between the software and all-optical ESNs arises primarily from differences in their nonlinear activation functions, as all other components are accurately implemented in both systems. In the instance shown in Figure 5, we see an NMSE of 0.57 from the software ESN and 0.54 from the optical ESN. These are median values from multiple computations, and are quite close. This means that our optical ESN performs nearly as good as the theoretical software ideal for the same size of 10 nodes in this particular benchmark, which is a strong result.
@@
Figure 7: Comparison of the performance of the Mackey-Glass time series prediction test. Parameters: $N=10$, $gL=12$, $\mathrm{Pump/Input}=1500$
@@ IV.3 Mackey-Glass
Finally, we use the Mackey-Glass task to test the performance of the ESN against the software-based version. This test particularly focuses on the dynamic memory retention capability of the all-optical ESN. This test requires an approximation of the Mackey-Glass chaotic time series equation [35] given by
@@
The values chosen for the constants in Equation (11) are the standard values $(\beta,\tau,n,\gamma)=(0.2,17,10,0.1)$. The solution to this equation is numerically approximated and allowed to run for 1000 steps before being used for the task. The task given to the ESN is to predict 5 steps ahead of the input, i.e. given the input $u_{k}=y_{k-5}$} we want a successful prediction of $y_{k}$}. This benchmark evaluates the reservoir computer’s memory capacity and its ability to predict chaotic temporal variations. Strong performance in this test indicates that the optical reservoir has sufficient long-term memory for effective time-based predictions, as accurately forecasting future values of a complex analytical time-series requires robust extrapolation capabilities derived from effective dynamic memory retention. We evaluate the same benchmark in the software-based sigmoid ESN, and compare the two to estimate the theoretical fit of the optical ESN. We used the following parameters for the all-optical ESN: $N=10,gL=12,$ and a pump/input ratio of 1500.
@@
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.
@@ IV.4 Noise Performance and Negative Values
For the case of a weak input probe undergoing high gain amplification, Brillouin amplification is known to be noisy [36]. This noise scales with pump intensity and is more prominent in cases with pump beam aberration, which can cause reduced performance by the optical reservoir [37]. In our design, these intense gain values are not necessary for the matrix multiplication steps, making our reservoir relatively safe from high gain noise. To further reduce the impact, the amplified spontaneous noise (ASE) can be mitigated by the use of a band-pass filter [22] and in our simulations, we see good performance without requiring output probe powers over 31 dBm, which would result in a negligible noise level.
@@
Another source of noise is optical power detection noise. This is accounted for by a simple gaussian noise model. Figure 9 shows examples of the performance of the optical ESN with detector noise included, compared with Figures 3, 5, and 7. For this test, we assumed the noise property of $0.32\%$ uncertainty at 15501550nm wavelength, which is a standard treatment of considering detector noise [38]. In the presence of noise, the nonlinear capability is severely reduced as can be seen in the sine-square classification test results. This is mitigated by choosing an optimal $gL$ value with a large pump splitting ratio, as detailed in the previous section. We observe from Figure 10 that the average NMSE begins to trend upward, unlike the noise free cases of Figures 4, 6, and 8. The cause of this is likely that the saturation point for the specific task as a function of pump/input ratio is reached earlier because of the presence of detection noise.
The optical ESN shows robustness in its time-series prediction ability, even in the presence of noise. The reported median NMSE in Figure 10 for the Mackey-Glass test is 0.15. This shows that the long-term memory capability is not diminished by the presence of noise, although there is a plateau in performance with varying pump splitting ratio, as can be seen from Figure 10 (c). We can see that overall, after accounting for power detection noise in the fiber setup, there is a reduction in performance of around 0.10.50.1-0.5 NMSE with more reduction in the classification task and less effect on prediction, implying that the efficiency of nonlinearity is affected more by the noise than the long and short term memory capabilities of the optical reservoir. These noise properties are for commercially available optical detectors. Hence, the negative impact of noise on performance can be further mitigated with the use of low noise detectors [39], [40].
One drawback of the selected benchmark tests is that they do not cover negative values. In order to demonstrate the ability of our design to access and predict negative inputs and targets, we use the Lorenz system test. It involves predicting the x-coordinate of a chaotic time series function [41]. The results for this test are shown in Figure 11.
@@ V Conclusion
We have provided the design for a feasible all-optical ESN with minimal resource overhead and energy requirements. For the nonlinear activation and matrix multiplication, we use SBS in two regimes: the large pump regime for linear behavior, and the self-pumped regime for nonlinear behavior. In the linear regime, tuning the pump power along with required attenuation using VOAs is enough to achieve the necessary multiplications. For the nonlinear activation step, according to reported achievable values of Brillouin gain in highly nonlinear fibers (HNLFs) [22], our design requires a fiber of length 363-6 m to achieve the required nonlinearity for good results, which is a very practical length of fiber. Such fibers also have a reasonable threshold power to show the required nonlinear output.
Additionally, our design requires only a single Brillouin fiber amplifier as a nonlinear activation device, making it highly scalable. However, a trade-off in computation speed has been introduced to allow for this scalable design. It is possible to achieve faster operation speeds at the cost of scalability, but this also greatly increases the hardware resource requirement, so we have decided against this approach. The nonlinear activation is optimizable by tuning the $gL$ value of the SBS amplifier as well as the pump splitting ratio, which refers to the ratio of the pump to the input given to the self-pumped Brillouin amplifier. By doing so, we analyze the behavior of the optical ESN for multiple benchmark tests to find the most optimal activation function design, while ensuring low overall energy consumption.
Our matrix multiplication and beam combination steps make use of standard fiber optical components such as VOAs, fiber beam combiners, 1:N splitters, combiners, and a 1:N switch. We see comparable results to a standard software ESN using a sigmoid activation function. To show this, we chose three tasks: classification of sine and square waves, prediction of the NARMA10 sequence, and prediction of the Mackey-Glass chaotic time series. Each benchmark task demonstrates the nonlinearity, polynomial algebra capability, and long-term memory of this ESN respectively, showing that universal computations are indeed accessible with this hardware. We validated our numerical simulations of the designed all-optical ESNs by comparing them to a software ESN with a sigmoid activation function. The results demonstrated that both systems achieve similar performance when using an equal number of nodes. This shows the feasibility and effectiveness of an all-optical approach to building an ESN.

### v3

@@ IV Results and Discussion
Figure 3: Comparison of the performance of the Sine-Square Classification test. Parameters: $N=10$, $gL=12$, $m=1500$
Good performance in machine learning hinges on the ability to approximate nonlinear input–output relationships and capture complex temporal dynamics. According to the Stone–Weierstrass theorem [stone1948generalized], the polynomial algebra property is central to achieving universality. Therefore, we selected three distinct tasks to evaluate these critical capabilities. They are the sine-square classification test, prediction of the NARMA10 system, and the Mackey-Glass time series prediction. Each task evaluates mainly the reservoir computer nonlinearity, polynomial algebra capabilities, and dynamic memory retention, respectively. Additionally, we address the nonlinear properties of the SBS interaction and, thus, the optimization of the activation function. To achieve good results in the tasks mentioned above, we require strong enough nonlinearity, while maintaining the convergence conditions of the ESN. We vary the value of the $gL$ product as well as the pump splitting ratio $m$ with these constraints in mind to optimize the performance of the ESN in the selected benchmark tests.
@@
For each test, we analyze the effect of the nonlinear strength of the SBS interaction, given by the $gL$ product value of the medium and the pump splitting ratio $m$ for the self-pumped nonlinear activation element, which is given by Equation (8). Here, $g$ represents the Brillouin gain and $L$ represents the interaction length. These values appear in Equation (7). The $gL$ product decides the nonlinear shape and the parking point of the activation function, as shown in Figure 1. The pump splitting ratio independently changes the parking point of the activation function, which allows us to optimize it for greater performance. It is important to note however that increasing the splitting ratio increases the power budget of the system. The parking point of the tunable activation function moves between -5 dBm to -10 dBm for the range of gL values that we have used for testing in this work, as shown in Figure 1. This means that, around the parking points, the pump power will be 100300100-300 mW for a splitting ratio of 1000. Since we have allowed for a maximum output $x_{k}$ value of $31$ dBm, the maximum pump power required in this scenario is $3.2$ W, which is quite a feasible power budget for the nonlinear activation elements. Better performance can generally be achieved for larger splitting ratios, which means that the best performance is restricted by the energy availability.
@@
(c) Average NMSE vs. gLgL values for various NN, $m=1000$.
@@ IV.1 Sine Square Classification
This test is a time series classification task, commonly used in estimating the nonlinearity of computing reservoirs [dudas2023quantum]. The input given is a series of randomly arranged sine and square waves, each 8 steps long. The task given to the ESN is to classify the current wave as sine or square. This test helps us understand the nonlinearity of the ESN. Distinguishing between smoothly varying sine waves and abruptly changing square waves requires mapping the data into a higher-dimensional hyperspace (i.e., nonlinear transformation), making classification easier. The results are shown in Figure 3.
On average, we see similar performance in terms of the NMSE with the software ESN using a sigmoid activation performing marginally better. We emphasize that the comparison between the software ESN and the optical ESN is not intended to show that one approach is superior to the other. Rather, the comparison serves as a validation check for the numerical simulations of our optical ESN. We anticipate that the performance of the software ESN and the optical ESN will be quite similar, as both implementations rely on the same underlying model, differing only slightly in their nonlinear activation functions. The reported values for this instance given in the figures are 0.14 for the software ESN and 0.12 for the optical ESN. These are median values from multiple computations.
@@
Figure 5: Comparison of the performance of the NARMA10 test. Parameters: $N=10$, $gL=12$, $m=1500$
The values reported by Slinkov et. al [slinkov2024all] allow the possibility of $gL$ values from 373-7 with fiber lengths of 363-6m at reasonable threshold powers of $1.26$ W which are very practical lengths of fiber. Experimental demonstrations have shown much higher levels of nonlinearity in the right materials [eggleton2019brillouin, choudhary2016advanced]. This shows promise for the translation of this architecture into an on-chip implementation, which has the benefit of improved scalability while maintaining the other advantages of this fiber-based approach.
@@ IV.2 NARMA10
(c) Average NMSE vs. gLgL values for various NN, $m=1000$.
@@
Next, the nonlinear autoregressive moving average (NARMA) [atiya2000new] task is used to test the ESN. This task mainly focuses on the polynomial algebra property of the optical ESN reservoir. The model for NARMA10, a version of NARMA popularly used for testing reservoir computers, is given by
@@
For this test, we used the following parameters of the optical ESN: $N=10,gL=12,$ and the pump splitting ratio of $m=1500$. On average we see marginally better performance from the software sigmoid ESN, but the NMSE values are comparable. We note that the performance difference between the software and optical ESNs arises primarily from differences in their nonlinear activation functions, as all other components are accurately implemented in both systems. In the instance shown in Figure 5, we see an NMSE of 0.57 from the software ESN and 0.54 from the optical ESN. These are median values from multiple computations, and are quite close. This means that our optical ESN performs nearly as good as the theoretical software ideal for the same size of 10 nodes in this particular benchmark, which is a strong result.
@@
Figure 7: Comparison of the performance of the Mackey-Glass time series prediction test. Parameters: $N=10$, $gL=12$, $m=1500$
@@ IV.3 Mackey-Glass
Finally, we use the Mackey-Glass task to test the performance of the ESN against the software-based version. This test particularly focuses on the dynamic memory retention capability of the optical ESN. This test requires an approximation of the Mackey-Glass chaotic time series equation [mackey1977oscillation] given by
@@
The values chosen for the constants in Equation (18) are the standard values $(\beta,\tau,n,\gamma)=(0.2,17,10,0.1)$. The solution to this equation is numerically approximated and allowed to run for 1000 steps before being used for the task. The task given to the ESN is to predict 5 steps ahead of the input, i.e. given the input $u_{k}=y_{k-5}$} we want a successful prediction of $y_{k}$}. This benchmark evaluates the reservoir computer’s memory capacity and its ability to predict chaotic temporal variations. Strong performance in this test indicates that the optical reservoir has sufficient long-term memory for effective time-based predictions, as accurately forecasting future values of a complex analytical time-series requires robust extrapolation capabilities derived from effective dynamic memory retention. We evaluate the same benchmark in the software-based sigmoid ESN, and compare the two to estimate the theoretical fit of the optical ESN. We used the following parameters for the optical ESN: $N=10,gL=12,$ and a pump splitting ratio of $m=1500$.
@@
(c) Average NMSE vs. gLgL values for various NN, $m=1000$.
@@ IV.4 Noise Performance and Negative Values
For the case of a weak input probe undergoing high gain amplification, Brillouin amplification is known to be noisy [scott1990gain]. This noise scales with pump intensity and is more prominent in cases with pump beam aberration, which can cause reduced performance by the optical reservoir [nathe2023reservoir]. In our design, these intense gain values are not necessary for the matrix multiplication steps, making our reservoir relatively safe from high gain noise. To further reduce the impact, the amplified spontaneous noise (ASE) can be mitigated by the use of a band-pass filter [slinkov2024all] and in our simulations, we see good performance without requiring output probe powers over 31 dBm, which would result in a negligible noise level.
@@
Another source of noise is optical power detection noise. This is accounted for by a simple gaussian noise model. Figure 9 shows examples of the performance of the optical ESN with detector noise included, compared with Figures 3, 5, and 7. For this test, we assumed the noise property of $0.32\%$ uncertainty at 15501550nm wavelength, which is a standard treatment of considering detector noise [vayshenker2000optical]. In the presence of noise, the nonlinear capability is severely reduced as can be seen in the sine-square classification test results. This is mitigated by choosing an optimal $gL$ value with a large pump splitting ratio, as detailed in the previous section. We observe from Figure 10 that the average NMSE begins to trend upward, unlike the noise free cases of Figures 4, 6, and 8. The cause of this is likely that the saturation point for the specific task as a function of the pump splitting ratio $m$ is reached earlier because of the presence of detection noise.
The optical ESN shows robustness in its time-series prediction ability, even in the presence of noise. The reported median NMSE in Figure 10 for the Mackey-Glass test is 0.15. This shows that the long-term memory capability is not diminished by the presence of noise, although there is a plateau in performance with varying pump splitting ratio, as can be seen from Figure 10 (c). We can see that overall, after accounting for power detection noise in the fiber setup, there is a reduction in performance of around 0.10.50.1-0.5 NMSE with more reduction in the classification task and less effect on prediction, implying that the efficiency of nonlinearity is affected more by the noise than the long and short term memory capabilities of the optical reservoir. These noise properties are for commercially available optical detectors. Hence, the negative impact of noise on performance can be further mitigated with the use of low noise detectors [lu2018ultra], [gasmi2016fast].
One drawback of the selected benchmark tests is that they do not cover negative values. In order to demonstrate the ability of our design to access and predict negative inputs and targets, we use the Lorenz system test. It involves predicting the x-coordinate of a chaotic time series function [lorenz1963deterministic]. The results for this test are shown in Figure 11.
@@ V Conclusion
We have provided the design for a feasible optical ESN with minimal resource overhead and energy requirements. For the nonlinear activation and matrix multiplication, we use SBS in two regimes: the large pump regime for linear behavior, and the self-pumped regime for nonlinear behavior. In the linear regime, tuning the pump power along with required attenuation using VOAs is enough to achieve the necessary multiplications and accumulations. For the nonlinear activation step, according to reported achievable values of Brillouin gain in highly nonlinear fibers (HNLFs) [slinkov2024all], our design requires a fiber of length 363-6 m to achieve the required nonlinearity for good results, which is a very practical length of fiber. Such fibers also have a reasonable threshold power to show the required nonlinear output.
Additionally, our design scales linearly with the number of nodes $N$, making it highly scalable compared to a standard matrix implementation that scales as $N^{2}$. It also uses only a single nonlinear activation element. However, a trade-off in computation speed has been introduced to allow for this scalable design. It is possible to achieve faster operation speeds at the cost of scalability, but this also greatly increases the hardware resource requirement, so we have decided against this approach. The nonlinear activation is optimizable by tuning the $gL$ value of the SBS amplifier as well as the pump splitting ratio $m$, which refers to the ratio of the pump to the input given to the self-pumped Brillouin amplifier. By doing so, we analyze the behavior of the optical ESN for multiple benchmark tests to find the most optimal activation function design, while ensuring low overall energy consumption.
Our matrix multiplication and accumulation steps make use of standard fiber optical components such as VOAs and fiber switches. We see comparable results to a standard software ESN using a sigmoid activation function. To show this, we chose three tasks: classification of sine and square waves, prediction of the NARMA10 sequence, and prediction of the Mackey-Glass chaotic time series. Each benchmark task demonstrates the nonlinearity, polynomial algebra capability, and long-term memory of this ESN respectively, showing that universal computations are indeed accessible with this hardware. We validated our numerical simulations of the designed optical ESNs by comparing them to a software ESN with a sigmoid activation function. The results demonstrated that both systems achieve similar performance when using an equal number of nodes. This shows the feasibility and effectiveness of an optical approach to building an ESN.